### v1

# MERLIN: Multimodal Embedding Refinement via LLM-based Iterative Navigation for Text-Video Retrieval-Rerank Pipeline

Abstract
The rapid expansion of multimedia content has made accurately retrieving relevant videos from large collections increasingly challenging. Recent advancements in text-video retrieval have focused on cross-modal interactions, large-scale foundation model training, and probabilistic modeling, yet often neglect the crucial user perspective, leading to discrepancies between user queries and the content retrieved. To address this, we introduce MERLIN (Multimodal Embedding Refinement via LLM-based Iterative Navigation), a novel, training-free pipeline that leverages Large Language Models (LLMs) for iterative feedback learning. MERLIN refines query embeddings from a user perspective, enhancing alignment between queries and video content through a dynamic question answering process. Experimental results on datasets like MSR-VTT, MSVD, and ActivityNet demonstrate that MERLIN substantially improves Recall@1, outperforming existing systems and confirming the benefits of integrating LLMs into multimodal retrieval systems for more responsive and context-aware multimedia retrieval [^1] Code will be released upon acceptance..
MERLIN: Multimodal Embedding Refinement via LLM-based Iterative Navigation for Text-Video Retrieval-Rerank Pipeline
Donghoon Han1111, Eunhwan Park1111, Gisang Lee2111, Adam Lee3, 4, Nojun Kwak5222 1Buzzni AI Lab, 2KAIST, 3UC Berkeley, 4Wrtn, 5Seoul National University {owen, jude}@buzzni.com, bobopack@kaist.ac.kr, alee00@berkeley.edu, nojunk@snu.ac.kr
footnotetext: 111 Equal contribution.footnotetext: 222 Corresponding author.
1 Introduction
Multimedia content has recently grown rapidly in both quantity and quality, making the task of finding relevant videos from vast collections increasingly challenging. While recent studies on text-video retrieval have primarily focused on cross-modal interaction Wang et al. (2023); Huang et al. (2023); Wu et al. (2023); Jin et al. (2023), large-scale foundation model training Chen et al. (2024b, 2023); Zhao et al. (2024); Wang et al. (2024a) and probabilistic modeling Hao et al. (2023); Fang et al. (2023); Hao and Zhang (2023), there remains a notable lack of consideration for the discrepancy in text-video retrieval. For instance, as illustrated in Figure 1, the video caption “a baby playing with a cat’s tail” fails to fully capture the additional context of a playful interaction between the baby and the cat. In real-world scenarios, such discrepancies often arise because users tend to submit succinct queries that do not capture the full context of the videos related to their search intent. Consequently, this mismatch can lead to unsatisfactory retrieval performance. Moreover, neglecting the user perspective, making users to refine their natural language query for multiple times to fully reflect their search intent, degrades the quality of user experience and makes it difficult to understand the search intent, leading to a discrepancy between user queries and the information within the retrieved videos.
Figure 1: An illustration of the discrepancy between the video caption which could be treated as a user query and the video from MSR-VTT dataset. blue and red indicate the additional details within the video frame, and those reflecting temporal information, respectively.
To address this issue, we introduce MERLIN (Multimodal Embedding Refinement via LLM-based Iterative Navigation), a novel training-free and iterative feedback learning pipeline that leverages the power of Large Language Models (LLMs) to augment queries based on the user perspective, thereby mitigating the aforementioned discrepancies and significantly improving the text-video retrieval performance. Inspired by human problem-solving and cognitive feedback mechanisms Flower and Hayes (1981); Doherty and Balzer (1988), we employ an interactive and iterative feedback learning Böhm et al. (2019); Stiennon et al. (2020); Ziegler et al. (2019); Wu et al. (2020); Ouyang et al. (2022); Glaese et al. (2022); Akyürek et al. (2023); Madaan et al. (2023); Lee et al. (2024); Liang et al. (2024) consisting of a question-and-answer process that iteratively refines query embeddings for text-video retrieval. Moreover, to our best knowledge, MERLIN presents the first implementation of a retrieval–rerank pipeline in the domain of text-video retrieval, establishing a novel framework that prioritizes user intention and interaction in refining search results.
The primary strength of MERLIN lies in its capability to iteratively adapt and refine query embeddings without necessitating the costly re-training of pre-trained models. By generating a series of questions based on the metadata and content of the top-retrieved video and answering the question through LLMs, MERLIN iteratively refines query embedding with answer embedding, thereby enhancing semantic relevance of queries and the information contained within the videos. Experimental results on benchmark datasets, including MSR-VTT, MSVD, and ActivityNet, demonstrate the superiority of the retrieval performance (e.g. Recall@K) by showing significant improvement. Specifically, MERLIN boosts text-video retrieval performance (Recall@1) of Google Multimodal Embedding from $44.00$ to $78.00$ on MSR-VTT, from $52.39$ to $77.61$ on MSVD and from $56.58$ to $68.44$ on ActivityNet.
The key contributions of our paper are as follows: (1) Introduction of MERLIN, a novel LLM-based framework for multimodal embedding refinement that addresses discrepancies between user queries and video content by integrating user perspectives. (2) Implementation of an iterative, cost-effective method for refining query embeddings using LLMs, significantly reducing computational demands while improving retrieval accuracy. (3) Presentation of the first retrieval-rerank pipeline in text-video retrieval, enhancing interactivity and context-awareness within multimodal systems. (4) Experimental results shows that MERLIN substantially improves Recall@1 on MSR-VTT, MSVD and ActivityNet, thereby demonstrating notable enhancements in zero-shot text-video retrieval.
Figure 2: An illustration of MERLIN for text-video retrieval. The yellow arrow represents the LLM Questioner returning a question for next round based on metadata of anchor video (Section 3.3). The green arrow represents the human-simulating LLM agent returning an answer based on the “video in mind” through Question Answering module and Aggregation module (Section 3.4). The pink arrow represents MERLIN returning a retrieved video candidates through Multimodal Encoder and Reranker (Section 3.5). The system initially retrieves video candidates $\hat{v}^{0}$ based on the input query text $q$ using a pre-trained multimodal encoder. Using anchor video, LLM Question Generator produces a question $\hat{q}^{1}$ to elicit additional information from the user (Section 3.3). The LLM Agent answers this question based on the “video in mind”, mimicking the human feedback process $\tilde{a}^{1}$. The query and answer embeddings are then gradually integrated for each round. The updated query embedding is used to rerank the video candidates $\hat{v}^{1}$, and the process repeats for multiple rounds.
2 Related Works
Dataset.
Text-to-video retrieval aims to retrieve relevant videos based on natural language descriptions and several benchmark video datasets Anne Hendricks et al. (2017); Caba Heilbron et al. (2015); Chen and Dolan (2011); Xu et al. (2016) have been curated for this task. One notable dataset is ActivityNet Caba Heilbron et al. (2015), which consists of video-text pairs capturing various human activities. Another widely used dataset is MSR-VTT Xu et al. (2016), which comprises open-domain web videos paired with natural language descriptions. These datasets provide a diverse range of video content and textual queries, enabling comprehensive evaluation of retrieval systems.
Method.
Prior studies have focused on cross-modal interaction, large-scale foundation model training, and probabilistic modeling. In cross-modal interaction Wang et al. (2023); Huang et al. (2023); Jin et al. (2023) have enhanced reasoning abilities by capturing cross-modal similarities at multiple granularity levels, introduced efficient video prompt mechanisms Lester et al. (2021) with minimal trainable parameters, and improved retrieval with strategies like Disentangled Conceptualization and Set-to-Set Alignment. In foundation model training Chen et al. (2024b, 2023); Zhao et al. (2024); Wang et al. (2024a) significant advances have been made with the development of large-scale video and vision-language models leveraging extensive web data, and fine-tuning techniques for better performance on downstream tasks. In probabilistic modeling Hao et al. (2023); Fang et al. (2023); Hao and Zhang (2023), novel alignment methods and modeling of video and text representations as probabilistic distributions have been proposed to improve text-video retrieval accuracy and address domain adaptation challenges.
Concurrent to prior studies and our work, Levy et al. (2023) proposed a chat-based image retrieval system (ChatIR) that interacts with users through conversation to gather additional information beyond the initial query, aiming to better understand and clarify the user’s search intent. Different from ChatIR, MERLIN incorporates frame-level answer generation tailored to the specific requirements of text-video retrieval, employing a training-free approach. Furthermore, inspired by Composed Image Retrieval Liu et al. (2021); Jang et al. (2024), we iteratively refine the embedding by employing spherical linear interpolation, instead iteratively concatenate question and answer pair and feed into retrieval model. Lastly, we handle both multi-modality data simultaneously, meaning that our generation result would be more likely aligned to user’s search intent. This iterative refinement process mirrors human tendencies to continuously improve their queries based on interactive feedback, akin to strategies seen in feedback-based refinement in textual content. This approach is supported by the growing application of reinforcement learning, which has been increasingly utilized to enhance the quality of generated content through both reference-based and reference-independent feedback mechanisms Böhm et al. (2019); Stiennon et al. (2020); Ziegler et al. (2019); Wu et al. (2020); Ouyang et al. (2022); Glaese et al. (2022); Akyürek et al. (2023); Madaan et al. (2023); Lee et al. (2024); Liang et al. (2024).
3 Multimodal Embedding Refinement via LLM based Iterative Navigation
3.1 Retrieval Scenario
Figure 3: An article of scenario of Google Lens.
In this work, we consider a scenario where a user aims to find a specific video in a search engine, having a firm intent and precise specifications in mind for the desired video. We refer to this desired video as the “video in mind” – an imagined video that the user seeks to retrieve, encompassing various attributes such as visual content, context, and temporal dynamics. However, expressing the full specifications of this “video in mind” through a single query often leads to a discrepancy between the user’s intent and the retrieved results, hindering the search experience.
We incorporate this “video in mind” scenario into the text-video retrieval task, where the goal is to retrieve the most relevant video from a collection based on the user’s textual query. Inspired by the success of multimodal systems like Google Lens [^2] https://lens.google/, which effectively handle and integrate both image and text inputs, we propose MERLIN, an interactive pipeline that dynamically elicits and incorporates user feedback to bridge the gap between the user’s search intent and the retrieved video candidates. MERLIN leverages the power of LLMs to simulate a human-like feedback process, where a “human-simulating LLM agent” engages in a multi-turn conversation with the user, asking clarifying questions and providing answers based on the “video in mind” – the imagined video that the agent has in mind based on the user’s initial query and subsequent feedback.
3.2 Background
Suppose that we have the query text $q\in\mathcal{Q}$, a video $v\in\mathcal{V}$, where $\mathcal{Q}$ and $\mathcal{V}$ indicate a set of queries and videos. Using a pre-trained multimodal encoder $f_{\text{enc}}$, we obtain the query and video embeddings ($\mathbf{e}_{q},\mathbf{e}_{v})$ as follows:
$\mathbf{e}_{q}$ $=f_{\text{enc}}(q)\in\mathbb{R}^{d_{q}}$
$\mathbf{e}_{v}$ $=f_{\text{enc}}(v)\in\mathbb{R}^{d_{v}},$
where $d_{q}$ and $d_{v}$ denote the dimension of query and video embeddings, respectively. The goal of text-video retrieval is to search the most relevant videos $\hat{v}$’s from a collection of videos $\mathcal{V}$ given a query text $q$ as follows:
$\big{[}\hat{v}_{0},\dots,\hat{v}_{k-1}\big{]}=\textsc{top-k}_{v\in\mathcal{V}}\big{(}\textsc{sim}(\mathbf{e}_{q},\mathbf{e}_{v})\big{)},$ (1)
where $\textsc{sim}(\cdot)$ is a similarity function (e.g., cosine distance, etc). Additionally, our system utilizes two key components: $\mathcal{M}$ and $\mathcal{T}$. Here, $\mathcal{M}$ represents the LLMs and template function $\mathcal{T}$ applies a pre-defined template to inputs [^3] Note that $\mathcal{M}$ is used interchangeably to indicate both a Large Language Model (LLM) and a Multimodal Large Language Model (MLLM)., [^4] Here, subscripts have been omitted for simplicity. However, subscripts are employed in the equations for each specific module (e.g., $\mathcal{M}_{\mathsf{question}}$). In addition, the pre-defined template is presented in Appendix due to the limited space.. Based on this background, we would like to introduce LLM-based iterative navigation, involving multiple rounds of feedback learning and reranking, leading to better performance and interpretability.
3.3 Question Generation
Suppose that we have retrieved candidates $\hat{v}^{r}_{k}$ where $r$ and $k$ indicate the round and the index of the retrieved top $K$ candidates, respectively. We choose $\hat{v}^{r}_{0}$ as an anchor candidate and generate the question with $\mathcal{M}_{\mathsf{question}}$ as follows [^5] Note that we use the caption from metadata of $\hat{v}^{r}_{0}$ and assume that each video consists of $N$ frames.:
$\hat{q}^{r}=\mathcal{M}_{\mathsf{question}}\big{(}\mathcal{T}_{\mathsf{question}}(\hat{v}^{r}_{0})\big{)}$ (2)
Intuitively, top-ranked candidate is more likely to align with the user’s query. This implies that assessing retrieved candidates with question generated from $\hat{v}^{r}_{0}$ using LLMs would enhance retrieval performance and interpretability.
3.4 Human-Simulating Agent
Video Question Answering.
Our underlying assumption is mitigating the discrepancy between user queries and the information within the videos would be helpful for the better retrieval performance. To this end, we answer the question $\hat{q}^{r}$ with video $\hat{v}^{r}$ which consists of $N$ frames sampled per second as follows:
$\big{[}\hat{a}^{(r,0)},\dots,\hat{a}^{(r,N)}\big{]}=\mathcal{M}_{\mathsf{answer}}\big{(}\mathcal{T}_{\mathsf{answer}}(\hat{q}^{r}),\hat{v}^{r}_{0}\big{)}$ (3)
It is worth noting that in a real-world scenario, $\mathcal{M}_{\mathsf{answer}}$ could be replaced by a human to consider the user perspective better. Additionally, using $N$ frames allows us to efficiently handle the temporal information inherent to video, capturing the dynamic aspects of the content. This approach enhances our ability to provide a more comprehensive understanding and alignment with the user’s query.
Aggregation.
The individual generated answers for each frame $\big{[}\hat{a}^{(r,0)},\dots,\hat{a}^{(r,N)}\big{]}$ are now subsequently fed into an Aggregation Module which is designed to summarize the multiple frame-level answers into a coherent and concise response to the original query as follows:
$\tilde{a}^{r}=\mathcal{M}_{\mathsf{aggr}}\Big{(}\mathcal{T}_{\mathsf{aggr}}\big{(}[\hat{a}^{(r,0)},\dots,\hat{a}^{(r,N)}]\big{)}\Big{)}$ (4)
[…]
3.5 Iterative Embedding Refinement for Reranking
Initially, we obtain the answer embedding $\mathbf{e}^{\tilde{a}^{r}}$ using the multimodal encoder $f_{\text{enc}}$ as follows: $\mathbf{e}^{\tilde{a}^{r}}=f_{\mathsf{enc}}(\tilde{a}^{r})$. Our objective is to dynamically refine the embedding by combining the information from the current round’s answer with the previous round’s refined embedding. To this end, in the pursuit of refining embeddings iteratively to enhance retrieval performance, we employ a spherical linear interpolation (SLERP) Shoemake (1985), which is particularly appropriated for interpolating between embeddings on the unit sphere, preserving the norm and the geometric properties of the embeddings.
Given the embeddings $\mathbf{e}^{r-1}$ from the previous round and $\mathbf{e}^{\tilde{a}^{r}}$, the angle $\theta$ between them is computed as:
$\theta=\arccos(\mathbf{e}^{\tilde{a}^{r}}\cdot\mathbf{e}^{r-1})$ (5)
Note that the angle is essential for determining the interpolation path. Finally, the refined embedding for the current round $\mathbf{e}^{r}$ is then calculated as:
$\mathbf{e}^{r}=\frac{\sin((1-\alpha)\theta)}{\sin(\theta)}\cdot\mathbf{e}^{\tilde{a}^{r}}+\frac{\sin(\alpha\theta)}{\sin(\theta)}\cdot\mathbf{e}^{r-1},$ (6)
where $\alpha\in[0,1]$ is a hyperparameter that balances the influence of the current answer embedding and the previous refined embedding. This interpolation not only ensures a smooth transition across embedding spaces but also incorporates both the originality of the current response and the semantic context retained from prior interactions. We assume that the potential risk of the iterative embedding refinement is query drift Mitra et al. (1998); Zighelnic and Kurland (2008); Shtok et al. (2012), a common phenomenon in information retrieval where the focus inadvertently shifts away from the original query intent due to the inclusion of progressively accumulated details. To mitigate the potential risk, we set the $\alpha=0.8$, prioritizing the query and earlier answer embeddings over the most recent answers. We expect that this simple yet effective strategy would preserve the thematic integrity of the initial query, akin to human conversational patterns where early-mentioned topics typically set the context for the entire conversation.
Model Rounds MSR-VTT MSVD ActivityNet
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
VAST Chen et al. (2024a) - 49.30 68.30 73.90 - - - - - -
InternVideo2-6B Wang et al. (2024b) - 55.9 78.3 85.1 59.3 84.4 89.6 63.2 85.6 92.5
LanguageBind-H Zhu et al. (2023) - 44.8 70.0 78.7 53.9 80.4 87.8 41.0 68.4 80.8
VideoPrism-g Madan et al. (2024) - 39.7 63.7 - - - - 52.7 79.4 -
Marengo-2.6 Labs (2024) - 49.35 73.47 - - - - 55.36 82.55 -
MERLIN 0 44.40 67.60 76.20 52.39 77.16 84.78 56.58 84.77 91.73
1 56.40 80.00 87.00 61.94 85.97 91.79 59.96 89.01 93.91
2 66.40 86.00 92.80 67.61 90.45 94.63 62.68 90.42 94.34
3 72.60 91.80 95.60 71.79 91.79 96.87 66.05 90.97 95.54
4 76.20 93.40 97.00 74.78 93.28 96.87 67.14 91.08 95.54
5 78.00 94.20 96.80 77.61 94.48 97.31 68.44 91.95 96.63
Table 1: The performance of zero-shot text-video retrieval on MSR-VTT, MSVD, and ActivityNet.
4 Experimental Results
4.1 Setting
To utilize mutlimodal encoders and LLMs without needing private GPUs, we use Google Multimodal Embedding API [^6] https://cloud.google.com/generative-ai-studio for encoding video and text, and the OpenAI GPT-4o API Achiam et al. (2023) [^7] https://chat.openai.com/ for generating questions and answers. These APIs offers comparable performance and reproducibility on benchmarks without private GPUs.
We evaluate MERLIN across three datasets: MSR-VTT, MSVD, and ActivityNet. For MSR-VTT, we sampled 500 videos from its 1,000-sample validation split. From MSVD and ActivityNet, we sampled all 670 and 919 videos from their respective test sets. For videos with multiple captions, we randomly selected one query per video.
4.2 Performance on Text-Video Retrieval
The performance of our system is presented in Table 1, demonstrating its efficacy through multiple rounds of feedback learning, reflecting the system’s ability to iteratively to refine and incorporate feedback. Particularly, MERLIN shows significant improvements with each round of feedback: On the MSR-VTT dataset, MERLIN shows improvements from a R@1 of $44.40$ to $78.00$, on the MSVD from $52.39$ to $77.61$, and on ActivityNet from $56.58$ to $68.44$ by the final round.
This highlights MERLIN’s capacity to adapt and enhance its response through iterative feedback learning. Despite the distinct challenges posed by each dataset, MERLIN significantly boosts its performance, thereby affirming the effectiveness of leveraging iterative feedback learning to enhance text-video retrieval task.
4.3 Average Ranking of QA Rounds
Figure 4: An illustration of the average ranking of target video for each dataset.
In addition to the retrieval performance presented in Table 1, the effectiveness of the iterative query enrichment is further highlighted by examining the average ranking of the target videos across question-answer rounds. This analysis is helpful for understanding how the process enhances the ranking of the target videos. As illustrated Figure 4, the average ranking of the target video consistently improves each consecutive round across all datasets. For instance, on the MSR-VTT dataset, the average ranking significantly improves from $18.57$ in round 00 to $2.5$ by final round. Similar improvements are observed on other datasets, with the average ranking on MSVD improving from $13.84$ to $2.4$, and on ActivityNet from $6$ to $2.6$. This demonstrates the consistent improvement, thereby confirming the effectiveness of MERLIN in reranking through iterative feedback learning.
4.4 The Effect of $\alpha$ in Iterative Embedding Refinement
As mentioned in Section 3.5, our assumption is mitigating query drift would preserve the thematic integrity of the initial query by assigning high $\alpha$ value, prioritizing the query and earlier answer embeddings over the most recent answers.
To validate our assumption in contrast to the experiment’s higher $\alpha=0.8$, we conduct additional experiments with assigning a reduced value $\alpha=0.2$, which allows us to observe the impact of shifting emphasis towards the latest answers. The results on the MSR-VTT and MSVD datasets show that setting a lower $\alpha$ initially improves retrieval performance in early rounds but leads to a decline after a few rounds, indicating potential query drift. Furthermore, the average ranking of the target video deteriorates in later rounds, suggesting the query representation has deviated from the user’s original intent.
Specifically, for MSR-VTT, MERLIN got 44.4/67.60/76.20 for recall@1/5/10 at round 00 respectively but ended up with 61.6/81.20/87.00 respectively at round $5$. For MSVD, MERLIN got 52.39/77.16/84.78 for recall@1/5/10 at round 00 respectively but ended up with 56.87/78.51/84.63 respectively at round $5$.
5 Conclusion
In conclusion, the MERLIN framework addresses a critical gap in the field of text-video retrieval by integrating the often-overlooked user perspective into the retrieval process. This integration is achieved through a novel, training-free pipeline that utilizes LLMs for iterative feedback learning, allowing for the dynamic refinement of query embeddings based on user interactions. MERLIN not only aligns more closely with user intent but also enhances the overall search experience by reducing discrepancies between user queries and retrieved video content.
The implementation of MERLIN shows a significant advancement in multimedia retrieval, introducing the first retrieval-rerank pipeline in this domain. By incorporating iterative feedback mechanisms inspired by human cognitive processes, MERLIN facilitates a more aligned and context-aware approach to text-video retrieval. Our experimental results demonstrate the effectiveness of this approach, with substantial improvements in retrieval performance observed across MSR-VTT, MSVD, and ActivityNet datasets.
Limitations
While our results are promising, we acknowledge that we cannot provide a comprehensive guide for adapting MERLIN to different settings, as we have not extensively explored the impact of changing various components. However, the core principle of integrating user feedback to iteratively refine the query embedding appears to be a robust approach, regardless of pipeline components, the specific domain or data modality. Future work could investigate the generalization of MERLIN to other multimedia retrieval tasks and explore the optimal configurations for different scenarios.
Other limitation of our approach lies in the use of a human-simulating LLM agent for answering questions based on static video frames. While this agent aims to mimic the human feedback process, it lacks the capability to grasp temporal information and attributes that require a high-level understanding of motion and dynamics. Since the LLM agent first generates answers based on static images and then aggregates them, it struggles to capture knowledge about direction, speed, and other temporal aspects present in the videos.
Moreover, as most pre-trained video encoders also have shortcomings in effectively modeling temporal capabilities Liu et al. (2024), our video encoder may be affected by this limitation as well. This creates a kind of chicken-and-egg problem, where video encoders can benefit from temporal-rich information only when they can understand temporal information effectively. Conversely, even if the video question-answering module (or similar counterparts) can handle temporal-rich information, if the video encoder does not possess the same capability, it may not benefit from this information. This temporal modeling challenge is a prevalent issue that the community needs to address collectively.
Acknowledgements
This work was partially supported by Deep Learning Research Group AttentionX.
Appendix A Algorithm
Algorithm 1 Iterative video reranking with question answering rounds
1:encoder $f_{\text{enc}}()$, user query $q\in\mathcal{Q}$, video $v\in\mathcal{V}$, total question-answer round $R$, retrieved top-$k$ videos at round $r$ $\hat{v}^{r}$, $i$-th candidate among top-$k$ videos at round $r$ $\hat{v}^{r}_{i}$
2:Encode $\mathbf{e}_{q}=f_{\text{enc}}(q)$ given user query $q$
3:Encode $\mathbf{e}_{v}=f_{\text{enc}}(v)$ given video $v$
4:Retrieve $\hat{v}^{0}=\textsc{top-k}_{v\in\mathcal{V}}\big{(}\textsc{sim}(\mathbf{e}_{q},\mathbf{e}_{v})\big{)}$ (Equation 1)
5:Initialize message list $m=[]$
6:for $r=1$ to $R$ do
7:     Append metadata of $\hat{v}^{r-1}_{0}$ to $m$
8:     Generate question $\hat{q}^{r}=\mathcal{M}_{\mathsf{question}}(m)$ (Equation 2)
9:     Append $\hat{q}^{r}$ to $m$
10:     Generate frame-level answers $\big{[}\hat{a}^{(r,0)},\dots,\hat{a}^{(r,N)}\big{]}=\mathcal{M}_{\mathsf{answer}}(\hat{q}^{r}:\text{video in mind})$ (Equation 3)
11:     Aggregate frame-level answers $\tilde{a}^{r}=\mathcal{M}_{\mathsf{aggr}}([\hat{a}^{(r,0)},\dots,\hat{a}^{(r,N)}\big{]})$ (Equation 4)
12:     Encode $\mathbf{e}_{A^{r}}=f_{\text{enc}}(A^{r})$
13:     Refine embedding $e=Refine(e_{q},\cdots,e_{A^{r}})$
14:     Retrieve $\hat{v}^{r}=\textsc{top-k}_{v\in\mathcal{V}}\big{(}\textsc{sim}(\mathbf{e},\mathbf{e}_{v})\big{)}$ (Equation 1)
15:end for
16:return Reranked retrieved videos $\hat{v}^{r}_{k}$
Appendix B Prompt Template
B.1 Prompt for Question Generation Module
In order to get useful information from user, it is critical to ask good question that could draw user intention. As depicted in Table 2, we set top $1$ ranked video as anchor video and prompted GPT-4o to refer to anchor video’s metadata. In our case, we used the video’s caption as metadata. However, we believe that questions could be more diverse if we could use other data such as Automatic Speech Recognition (ASR) captions, the characteristics of the video, and so on. As MERLIN proceeds with the chat with the user (a user-simulating agent), we stacked previous questions and answers and To encourage GPT-4 to generate diverse questions without repeating previous ones..
B.2 Prompts for Human-Simulating Agent
As human-simulating agent has two steps for answering to the question regarding “video in mind”, we have two different settings for each step. This method lacks in understanding direction, speed and other temporal knowledge as we dealt in the Limitation. However, we experimentally showed that our human-simulating agent is helpful in enriching information.
B.2.1 Prompt for Question-Answering Module
As depicted in Table 3, we sampled frame from video for every 1 second. Then we asynchronously inputed sampled frame and question from MERLIN. We prompted GPT-4o to answer in detail about fact and not just answer in “Yes” or “No”. However this question-answering module is the part that takes up large portion of API cost so video may be sample in wider stride to lower the API cost.
B.2.2 Prompt for Aggregation Module
As depicted in Table 4, we aggregate all the answers generated from the question-answering module. We prompted GPT-4o to aggregate multiple answers made with multiple frame at question-answering module and appended an aggregating example.
Appendix C Case Studies
The main objective of MERLIN is to improve the ranking of failure cases where the target video is not among the top-ranked candidates. At the same time, it is important to keep the success case to stay in the top-ranked candidates while MERLIN proceeds chatting with user. Retrieving the target video among the top-ranked candidates indicates that MERLIN consistently reflects user intention during the conversation. To qualitatively verify that MERLIN performs its tasks according to the aforementioned objectives, we reviewed several case studies. We focused on how MERLIN bring the rank of failure cases.
C.1 Case study for ActivityNet
As shown in the Figure 5, initial ranking of target video was $224$ using paired query from dataset. However as MERLIN augmented query using user’s response, rank boosted to $36\rightarrow 14\rightarrow 4\rightarrow 1$ as round proceeds. During the conversation, MERLIN was able to understand that the user was looking for a video about Christmas themes, featuring two people, and involving gift wrapping. It managed to rank the target video on top with augmented information.
C.2 Case study for MSVD
As shown in Figure 6, initial ranking of target video was $154$ using paired query from dataset. However as MERLIN augmented query using user’s response, rank boosted to $14\rightarrow 1\rightarrow 1\rightarrow 1$ as round proceeds. During the conversation, MERLIN was able to understand that the user was looking for a video about NBA All-Star game, broadcasted on TNT and score board telling 74:75. It managed to rank the target video on top with augmented information at early round and manage to keep top rank during multiple rounds.
C.3 Case study for MSR-VTT
As shown in Figure 7, initial ranking of target video was $361$ using paired query from dataset. However as MERLIN augmented query using user’s response, rank boosted to $197\rightarrow 14\rightarrow 1\rightarrow 1$ as round proceeds. During the conversation, MERLIN was able to understand that the detailed feature and gesture of human featured on “video in mind”. It managed to rank the target video on top with augmented information at early round and manage to keep top rank during multiple rounds.
Details about question generation module in MERLIN
System prompt
You are given caption about certain video(anchor video) and query used to retrieve the anchor video. However this video may not be the exact video the I am looking for.
Your role is to ask question about the video I have in mind to get more information about video. You have 5 rounds and you can only ask one question at a time.
Focus on attributes like number of people, color, shape and etc.
Initial prompt
This is caption of retrieved video. Read the video captions and ask some question to gain more information to help find out exact video. Some video may not have caption due to API error saying sorry I can’t provide blah blah. Captions for video: {anchor video’s caption}
Question:
Question-answering round prompt
answer: {Aggregated answer from user-simulating Agent}
Based on answer, here’s caption of reranked video.
caption: {reranked top1 video caption as anchor caption}
Keep asking.
Question:
Max tokens
- 1500
Temperature
- 0.75
Table 2: The instruction and specification for question generation module in MERLIN using GPT-4o. After initial retrieval at round 00, MERLIN generates question with initial prompt using information of anchor video’s caption. After user answers to the question, MERLIN rerank the and generate question using new anchor and question-answering round prompt.
Details about human-simulating agent (question-answering module)
System Prompt
You are a helpful assistant that answer the question with details. Don’t jsut answer in yes or no. Provide more details(about facts) about the image that might help the questioner.
Input format
- text: {Question from MERLIN}
- image: {Image encoded in base64 captured from video in mind in 1 second interval.}
Max tokens
- 50
Temperature
- 0.3
Image sampling rate
- 1 second
Table 3: The instruction and specification for video question-answering human-simulating agent using GPT-4o (question-answering module).
Details about human-simulating agent (aggregation module)
System Prompt
The VQA model is designed to answer questions based on images. To apply it to videos, frames are uniformly extracted from the video over time, and the model provides an answer for each frame to a given question. This means that for a single question, there will be multiple answers - one for each extracted frame. Your role is to review all of the individual answers and summarize them to provide a final answer to the original question. When making final answer, don’t user unnecessary words like ‘Based on the individual answers provided by the VQA model,’. Just answer to the question.
For example, if the question is “Did a cookie appear in the video?” and the individual answers from the frames are [“No”, “No”, “Yes”, “No”], then since a cookie appeared in the 3rd frame, you should summarize and answer the question as “Yes”. Length of aggregated answer should be around 30~35 words.
Input format
Question: {Question from MERLIN}
VQA Answer: {Answers from question-answering module}
Aggregated Answer:
Max tokens
- 100
Temperature
- 0.5
Table 4: The instruction and specification for video question-answering human-simulating agent using GPT-4o (aggregation module).
Figure 5: Qualitative evaluation of MERLIN on ActivityNet. sample: v_juiMCvZUYwk.
Figure 6: Qualitative evaluation of MERLIN on MSVD. sample: hbE29pZh76I_3_8.
Figure 7: Qualitative evaluation of MERLIN on MSR-VTT1ka. sample: video8471.

### v2

@@
 Abstract
-The rapid expansion of multimedia content has made accurately retrieving relevant videos from large collections increasingly challenging.
+The rapid expansion of multimedia content has made it increasingly challenging to retrieve relevant videos from large collections accurately.
 Recent advancements in text-video retrieval have focused on cross-modal interactions, large-scale foundation model training, and probabilistic modeling, yet often neglect the crucial user perspective, leading to discrepancies between user queries and the content retrieved.
-To address this, we introduce MERLIN (Multimodal Embedding Refinement via LLM-based Iterative Navigation), a novel, training-free pipeline that leverages Large Language Models (LLMs) for iterative feedback learning.
+To address this, we introduce MERLIN (Multimodal Embedding Refinement via LLM-based Iterative Navigation), a novel training-free pipeline that leverages Large Language Models (LLMs) for iterative feedback learning.
 MERLIN refines query embeddings from a user perspective, enhancing alignment between queries and video content through a dynamic question answering process.
-Experimental results on datasets like MSR-VTT, MSVD, and ActivityNet demonstrate that MERLIN substantially improves Recall@1, outperforming existing systems and confirming the benefits of integrating LLMs into multimodal retrieval systems for more responsive and context-aware multimedia retrieval [^1] Code will be released upon acceptance..
+Experimental results on datasets like MSR-VTT, MSVD, and ActivityNet demonstrate that MERLIN substantially improves R@1, outperforming existing systems and confirming the benefits of integrating LLMs into multimodal retrieval systems for more responsive and context-aware multimedia retrieval [^1] https://github.com/dhk1349/MERLIN_text_to_video_search.git.
 MERLIN: Multimodal Embedding Refinement via LLM-based Iterative Navigation for Text-Video Retrieval-Rerank Pipeline
-Donghoon Han1111, Eunhwan Park1111, Gisang Lee2111, Adam Lee3, 4, Nojun Kwak5222 1Buzzni AI Lab, 2KAIST, 3UC Berkeley, 4Wrtn, 5Seoul National University {owen, jude}@buzzni.com, bobopack@kaist.ac.kr, alee00@berkeley.edu, nojunk@snu.ac.kr
-footnotetext: 111 Equal contribution.footnotetext: 222 Corresponding author.
+Donghoon Han1111, Eunhwan Park1111, Gisang Lee2, 3111, Adam Lee4, 5, Nojun Kwak6222 1Buzzni AI Lab, 2KAIST, 3Mathpresso Inc., 4UC Berkeley, 5Fainders AI, 6Seoul National University {owen, jude}@buzzni.com, bobopack@kaist.ac.kr, alee00@berkeley.edu, nojunk@snu.ac.kr
 1 Introduction
@@
 Consequently, this mismatch can lead to unsatisfactory retrieval performance.
-Moreover, neglecting the user perspective, making users to refine their natural language query for multiple times to fully reflect their search intent, degrades the quality of user experience and makes it difficult to understand the search intent, leading to a discrepancy between user queries and the information within the retrieved videos.
+Moreover, neglecting the user perspective makes users refine their natural language query multiple times to fully reflect their search intent.
+This degrades the quality of user experience and makes it difficult to understand the search intent, leading to a discrepancy between user queries and the information within the retrieved videos.
 Figure 1: An illustration of the discrepancy between the video caption which could be treated as a user query and the video from MSR-VTT dataset.
-blue and red indicate the additional details within the video frame, and those reflecting temporal information, respectively.
+Blue indicates the details that can be observed statically within the video frame, while red reflects the information that can be obtained temporally across multiple frames.
 To address this issue, we introduce MERLIN (Multimodal Embedding Refinement via LLM-based Iterative Navigation), a novel training-free and iterative feedback learning pipeline that leverages the power of Large Language Models (LLMs) to augment queries based on the user perspective, thereby mitigating the aforementioned discrepancies and significantly improving the text-video retrieval performance.
@@
 (2023); Lee et al.
-(2024); Liang et al.
-(2024) consisting of a question-and-answer process that iteratively refines query embeddings for text-video retrieval.
+(2024a); Liang et al.
+(2024) consisting of a question answering process that iteratively refines query embeddings for text-video retrieval.
 Moreover, to our best knowledge, MERLIN presents the first implementation of a retrieval–rerank pipeline in the domain of text-video retrieval, establishing a novel framework that prioritizes user intention and interaction in refining search results.
 The primary strength of MERLIN lies in its capability to iteratively adapt and refine query embeddings without necessitating the costly re-training of pre-trained models.
-By generating a series of questions based on the metadata and content of the top-retrieved video and answering the question through LLMs, MERLIN iteratively refines query embedding with answer embedding, thereby enhancing semantic relevance of queries and the information contained within the videos.
+As shown in Figure 2, when a user submits a query, MERLIN generates questions based on the metadata of the retrieved video candidates and presents these questions to the user.
+By gathering additional information from the user’s responses, MERLIN refines the embeddings to improve retrieval accuracy, thereby helping users find “video in mind”222“video in mind” refers to the specific video users are looking for or have in mind during the search process..
 Experimental results on benchmark datasets, including MSR-VTT, MSVD, and ActivityNet, demonstrate the superiority of the retrieval performance (e.g.
-Recall@K) by showing significant improvement.
-Specifically, MERLIN boosts text-video retrieval performance (Recall@1) of Google Multimodal Embedding from $44.00$ to $78.00$ on MSR-VTT, from $52.39$ to $77.61$ on MSVD and from $56.58$ to $68.44$ on ActivityNet.
+R@K) by showing significant improvement.
+Specifically, MERLIN boosts text-video retrieval performance (R@1) of Google Multimodal Embedding from $44.00$ to $78.00$ on MSR-VTT, from $52.39$ to $77.61$ on MSVD and from $56.58$ to $68.44$ on ActivityNet.
 The key contributions of our paper are as follows: (1) Introduction of MERLIN, a novel LLM-based framework for multimodal embedding refinement that addresses discrepancies between user queries and video content by integrating user perspectives.
@@
 (3) Presentation of the first retrieval-rerank pipeline in text-video retrieval, enhancing interactivity and context-awareness within multimodal systems.
-(4) Experimental results shows that MERLIN substantially improves Recall@1 on MSR-VTT, MSVD and ActivityNet, thereby demonstrating notable enhancements in zero-shot text-video retrieval.
+(4) Experimental results shows that MERLIN substantially improves R@1 on MSR-VTT, MSVD and ActivityNet, thereby demonstrating notable enhancements in zero-shot text-video retrieval.
 Figure 2: An illustration of MERLIN for text-video retrieval.
@@
 The system initially retrieves video candidates $\hat{v}^{0}$ based on the input query text $q$ using a pre-trained multimodal encoder.
-Using anchor video, LLM Question Generator produces a question $\hat{q}^{1}$ to elicit additional information from the user (Section 3.3).
+Using this anchor video, LLM Question Generator produces a question $\hat{q}^{1}$ to elicit additional information from the user (Section 3.2).
 The LLM Agent answers this question based on the “video in mind”, mimicking the human feedback process $\tilde{a}^{1}$.
@@
 (2024); Wang et al.
-(2024a) significant advances have been made with the development of large-scale video and vision-language models leveraging extensive web data, and fine-tuning techniques for better performance on downstream tasks.
+(2024a), significant advances have been made with the development of large-scale video and vision-language models leveraging extensive web data, and fine-tuning techniques for better performance on downstream tasks.
 In probabilistic modeling Hao et al.
 (2023); Fang et al.
-(2023); Hao and Zhang (2023), novel alignment methods and modeling of video and text representations as probabilistic distributions have been proposed to improve text-video retrieval accuracy and address domain adaptation challenges.
-Concurrent to prior studies and our work, Levy et al.
+(2023); Hao and Zhang (2023), novel alignment methods and modeling of video and text representations as probabilistic distributions have been proposed to improve text-video retrieval accuracy and addressed domain adaptation challenges.
+Concurrent to prior studies, Levy et al.
 (2023) proposed a chat-based image retrieval system (ChatIR) that interacts with users through conversation to gather additional information beyond the initial query, aiming to better understand and clarify the user’s search intent.
-Different from ChatIR, MERLIN incorporates frame-level answer generation tailored to the specific requirements of text-video retrieval, employing a training-free approach.
+Following from ChatIR, Lee et al.
+(2024b) proposed the plug-and-play interactive text-to-image retrieval system.
+Different from ChatIR, our MERLIN incorporates frame-level answer generation tailored to the specific requirements of text-video retrieval, employing a training-free approach.
 Furthermore, inspired by Composed Image Retrieval Liu et al.
 (2021); Jang et al.
-(2024), we iteratively refine the embedding by employing spherical linear interpolation, instead iteratively concatenate question and answer pair and feed into retrieval model.
-Lastly, we handle both multi-modality data simultaneously, meaning that our generation result would be more likely aligned to user’s search intent.
+(2024), we iteratively refine the embedding by employing spherical linear interpolation, instead of iteratively concatenating question and answer pair and feeding into the retrieval model.
+Lastly, we handle both multi-modality data simultaneously, meaning that our generation result would be more likely aligned to the user’s search intent.
 This iterative refinement process mirrors human tendencies to continuously improve their queries based on interactive feedback, akin to strategies seen in feedback-based refinement in textual content.
@@
 (2023); Lee et al.
-(2024); Liang et al.
+(2024a); Liang et al.
 (2024).
 3 Multimodal Embedding Refinement via LLM based Iterative Navigation
-3.1 Retrieval Scenario
-Figure 3: An article of scenario of Google Lens.
-In this work, we consider a scenario where a user aims to find a specific video in a search engine, having a firm intent and precise specifications in mind for the desired video.
-We refer to this desired video as the “video in mind” – an imagined video that the user seeks to retrieve, encompassing various attributes such as visual content, context, and temporal dynamics.
-However, expressing the full specifications of this “video in mind” through a single query often leads to a discrepancy between the user’s intent and the retrieved results, hindering the search experience.
-We incorporate this “video in mind” scenario into the text-video retrieval task, where the goal is to retrieve the most relevant video from a collection based on the user’s textual query.
-Inspired by the success of multimodal systems like Google Lens [^2] https://lens.google/, which effectively handle and integrate both image and text inputs, we propose MERLIN, an interactive pipeline that dynamically elicits and incorporates user feedback to bridge the gap between the user’s search intent and the retrieved video candidates.
-MERLIN leverages the power of LLMs to simulate a human-like feedback process, where a “human-simulating LLM agent” engages in a multi-turn conversation with the user, asking clarifying questions and providing answers based on the “video in mind” – the imagined video that the agent has in mind based on the user’s initial query and subsequent feedback.
 3.1 Background
+Algorithm 1 Iterative video reranking with question answering rounds
+1:encoder $f_{\text{enc}}()$, user query $q\in\mathcal{Q}$, video $v\in\mathcal{V}$, total question answer round $R$, retrieved top-$k$ videos at round $r$ $\hat{v}^{r}$, $i$-th candidate among top-$k$ videos at round $r$ $\hat{v}^{r}_{i}$, $v^{m}$ a video that user is looking for
+2:Encode $\mathbf{e}_{q}=f_{\text{enc}}(q)$ given user query $q$
+3:Encode $\mathbf{e}_{v}=f_{\text{enc}}(v)$ given video $v$
+4:Retrieve $\hat{v}^{0}=\textsc{top-k}_{v\in\mathcal{V}}\big{(}\textsc{sim}(\mathbf{e}_{q},\mathbf{e}_{v})\big{)}$ (Equation 1)
+5:Initialize message list $m=[]$
+6:for $r=1$ to $R$ do
+7:     Append metadata of $\hat{v}^{r-1}_{0}$ to $m$
+8:     Generate question $\hat{q}^{r}=\mathcal{M}_{\mathsf{question}}(m)$ (Equation 2)
+9:     Append $\hat{q}^{r}$ to $m$
+10:     Generate frame-level answers $\big{[}\hat{a}^{(r,0)},\dots,\hat{a}^{(r,N)}\big{]}=\mathcal{M}_{\mathsf{answer}}(\hat{q}^{r}:v^{m})$ (Equation 3)
+11:     Aggregate frame-level answers $\tilde{a}^{r}=\mathcal{M}_{\mathsf{aggr}}([\hat{a}^{(r,0)},\dots,\hat{a}^{(r,N)}\big{]})$ (Equation 4)
+12:     Encode $\mathbf{e}_{A^{r}}=f_{\text{enc}}(A^{r})$
+13:     Refine embedding $e=\textsc{Refine}(e_{q},\cdots,e_{A^{r}})$ (Equation 6)
+14:     Retrieve $\hat{v}^{r}=\textsc{top-k}_{v\in\mathcal{V}}\big{(}\textsc{sim}(\mathbf{e},\mathbf{e}_{v})\big{)}$ (Equation 1)
+15:end for
+16:return Reranked retrieved videos $\hat{v}^{r}_{k}$
 Suppose that we have the query text $q\in\mathcal{Q}$, a video $v\in\mathcal{V}$, where $\mathcal{Q}$ and $\mathcal{V}$ indicate a set of queries and videos.
 Using a pre-trained multimodal encoder $f_{\text{enc}}$, we obtain the query and video embeddings ($\mathbf{e}_{q},\mathbf{e}_{v})$ as follows:
-$\mathbf{e}_{q}$ $=f_{\text{enc}}(q)\in\mathbb{R}^{d_{q}}$
-$\mathbf{e}_{v}$ $=f_{\text{enc}}(v)\in\mathbb{R}^{d_{v}},$
-where $d_{q}$ and $d_{v}$ denote the dimension of query and video embeddings, respectively.
+$\mathbf{e}_{q}$ $=f_{\text{enc}}(q)\in\mathbb{R}^{d}$
+$\mathbf{e}_{v}$ $=f_{\text{enc}}(v)\in\mathbb{R}^{d},$
+where $d$ denotes the dimension of embedding.
 The goal of text-video retrieval is to search the most relevant videos $\hat{v}$’s from a collection of videos $\mathcal{V}$ given a query text $q$ as follows:
@@
 Additionally, our system utilizes two key components: $\mathcal{M}$ and $\mathcal{T}$.
-Here, $\mathcal{M}$ represents the LLMs and template function $\mathcal{T}$ applies a pre-defined template to inputs [^3] Note that $\mathcal{M}$ is used interchangeably to indicate both a Large Language Model (LLM) and a Multimodal Large Language Model (MLLM)., [^4] Here, subscripts have been omitted for simplicity.
+Here, $\mathcal{M}$ represents the LLMs and template function $\mathcal{T}$ applies a pre-defined template to inputs [^3] Note that $\mathcal{M}$ is used interchangeably to indicate both a Large Language Model (LLM) and a Large Multimodal Model (LMM)., [^4] Here, subscripts have been omitted for simplicity.
 However, subscripts are employed in the equations for each specific module (e.g., $\mathcal{M}_{\mathsf{question}}$).
@@
 Suppose that we have retrieved candidates $\hat{v}^{r}_{k}$ where $r$ and $k$ indicate the round and the index of the retrieved top $K$ candidates, respectively.
-We choose $\hat{v}^{r}_{0}$ as an anchor candidate and generate the question with $\mathcal{M}_{\mathsf{question}}$ as follows [^5] Note that we use the caption from metadata of $\hat{v}^{r}_{0}$ and assume that each video consists of $N$ frames.:
-$\hat{q}^{r}=\mathcal{M}_{\mathsf{question}}\big{(}\mathcal{T}_{\mathsf{question}}(\hat{v}^{r}_{0})\big{)}$ (2)
+We choose $\hat{v}^{r-1}_{0}$ as an anchor candidate and generate the question with $\mathcal{M}_{\mathsf{question}}$ as follows [^5] Note that we use the caption from metadata of $\hat{v}^{r}_{0}$ and assume that each video consists of $N$ frames.:
+$\hat{q}^{r}=\mathcal{M}_{\mathsf{question}}\big{(}\mathcal{T}_{\mathsf{question}}(\hat{v}^{r-1}_{0})\big{)}.$ (2)
 Intuitively, top-ranked candidate is more likely to align with the user’s query.
-This implies that assessing retrieved candidates with question generated from $\hat{v}^{r}_{0}$ using LLMs would enhance retrieval performance and interpretability.
+This implies that assessing retrieved candidates with question generated from $\hat{v}^{r-1}_{0}$ using LLMs would enhance retrieval performance and interpretability.
 3.3 Human-Simulating Agent
@@
 Our underlying assumption is mitigating the discrepancy between user queries and the information within the videos would be helpful for the better retrieval performance.
-To this end, we answer the question $\hat{q}^{r}$ with video $\hat{v}^{r}$ which consists of $N$ frames sampled per second as follows:
-$\big{[}\hat{a}^{(r,0)},\dots,\hat{a}^{(r,N)}\big{]}=\mathcal{M}_{\mathsf{answer}}\big{(}\mathcal{T}_{\mathsf{answer}}(\hat{q}^{r}),\hat{v}^{r}_{0}\big{)}$ (3)
-It is worth noting that in a real-world scenario, $\mathcal{M}_{\mathsf{answer}}$ could be replaced by a human to consider the user perspective better.
+To this end, a human-simulating agent answers the question $\hat{q}^{r}$ with video in mind $v^{m}$, which consists of N frames sampled per second as follows.
+In this process, we assume a user searching for a specific video, and create a human-simulating agent to mimic the behavior of that user.
+The agent generates responses by referencing both the video in mind $v^{m}$ (the video the user is looking for) and the questions generated by MERLIN as following:
+$\big{[}\hat{a}^{(r,0)},\dots,\hat{a}^{(r,N)}\big{]}=\mathcal{M}_{\mathsf{answer}}\big{(}\mathcal{T}_{\mathsf{answer}}(\hat{q}^{r}),v^{m}\big{)}$ (3)
+It is worth noting that in a real-world scenario, $\mathcal{M}_{\mathsf{answer}}$ could be replaced by a human.
 Additionally, using $N$ frames allows us to efficiently handle the temporal information inherent to video, capturing the dynamic aspects of the content.
@@
 The individual generated answers for each frame $\big{[}\hat{a}^{(r,0)},\dots,\hat{a}^{(r,N)}\big{]}$ are now subsequently fed into an Aggregation Module which is designed to summarize the multiple frame-level answers into a coherent and concise response to the original query as follows:
-$\tilde{a}^{r}=\mathcal{M}_{\mathsf{aggr}}\Big{(}\mathcal{T}_{\mathsf{aggr}}\big{(}[\hat{a}^{(r,0)},\dots,\hat{a}^{(r,N)}]\big{)}\Big{)}$ (4)
+$\tilde{a}^{r}=\mathcal{M}_{\mathsf{aggr}}\Big{(}\mathcal{T}_{\mathsf{aggr}}\big{(}[\hat{a}^{(r,0)},\dots,\hat{a}^{(r,N)}]\big{)}\Big{)}.$ (4)
 […]
@@
 Given the embeddings $\mathbf{e}^{r-1}$ from the previous round and $\mathbf{e}^{\tilde{a}^{r}}$, the angle $\theta$ between them is computed as:
-$\theta=\arccos(\mathbf{e}^{\tilde{a}^{r}}\cdot\mathbf{e}^{r-1})$ (5)
+$\theta=\arccos(\mathbf{e}^{\tilde{a}^{r}}\cdot\mathbf{e}^{r-1}).$ (5)
 Note that the angle is essential for determining the interpolation path. Finally, the refined embedding for the current round $\mathbf{e}^{r}$ is then calculated as:
@@
 4.1 Setting
-To utilize mutlimodal encoders and LLMs without needing private GPUs, we use Google Multimodal Embedding API [^6] https://cloud.google.com/generative-ai-studio for encoding video and text, and the OpenAI GPT-4o API Achiam et al.
+To utilize multimodal encoders and LLMs without needing private GPUs, we use Google Multimodal Embedding API [^6] https://cloud.google.com/generative-ai-studio for encoding video and text, and the OpenAI GPT-4o API Achiam et al.
 (2023) [^7] https://chat.openai.com/ for generating questions and answers.
@@
 4.2 Performance on Text-Video Retrieval
-The performance of our system is presented in Table 1, demonstrating its efficacy through multiple rounds of feedback learning, reflecting the system’s ability to iteratively to refine and incorporate feedback.
-Particularly, MERLIN shows significant improvements with each round of feedback: On the MSR-VTT dataset, MERLIN shows improvements from a R@1 of $44.40$ to $78.00$, on the MSVD from $52.39$ to $77.61$, and on ActivityNet from $56.58$ to $68.44$ by the final round.
+The performance of our system is presented in Table 1, demonstrating its efficacy through multiple rounds of feedback learning, reflecting the system’s ability to iteratively refine and incorporate feedback.
+Particularly, MERLIN shows significant improvements with each round of feedback: On the MSR-VTT dataset, MERLIN shows improvements of R@1 from $44.40$ to $78.00$, on the MSVD from $52.39$ to $77.61$, and on ActivityNet from $56.58$ to $68.44$ by the final round.
 This highlights MERLIN’s capacity to adapt and enhance its response through iterative feedback learning.
@@
 Figure 3: An illustration of the average ranking of target video for each dataset.
-In addition to the retrieval performance presented in Table 1, the effectiveness of the iterative query enrichment is further highlighted by examining the average ranking of the target videos across question-answer rounds.
+In addition to the retrieval performance presented in Table 1, the effectiveness of the iterative query enrichment is further highlighted by examining the average ranking of the target videos across question answer rounds.
 This analysis is helpful for understanding how the process enhances the ranking of the target videos.
 As illustrated Figure 3, the average ranking of the target video consistently improves each consecutive round across all datasets.
-For instance, on the MSR-VTT dataset, the average ranking significantly improves from $18.57$ in round 00 to $2.5$ by final round.
+For instance, on the MSR-VTT dataset, the average ranking significantly improves from $18.57$ in round 00 to $2.5$ by the final round.
 Similar improvements are observed on other datasets, with the average ranking on MSVD improving from $13.84$ to $2.4$, and on ActivityNet from $6$ to $2.6$.
 This demonstrates the consistent improvement, thereby confirming the effectiveness of MERLIN in reranking through iterative feedback learning.
-4.4 The Effect of $\alpha$ in Iterative Embedding Refinement
+5 Ablation Study
+Model Rounds MSR-VTT
+R@1 R@5 R@10
+Final Query Retrieval (FQR) 5 51.40 71.00 78.80
+Refined Reranking (RR) 5 53.60 74.40 81.80
+MERLIN 5 78.00 94.20 96.80
+Table 2: The performance comparison of video retrieval performance on MSR-VTT using R@K between Final Query Retrieval (FQR), Refined Reranking (RR), and MERLIN.
+It is worth noting that FQR and RR employ the generated query at final round.
+The iterative embedding refinement improves retrieval performance.
+The results in Table 2 demonstrate the effectiveness of iterative embedding refinement in improving the retrieval performance.
+Final Query Retrieval (FQR), which direct retrieves videos using the generate query, achieves a R@1 of $51.40$.
+Refined Reranking (RR), which applies reranking to the top-100 initial results, improves performance to $53.60$ at R@1.
+However, MERLIN, which leverages iterative refinement through multiple rounds of interaction between the query and video embeddings, significantly outperforms both methods, reaching a R@1 of $78.00$, demonstrating the advantage of iterative refinement for aligning query representations with video content.
+The consistent improvements at R@5 and R@10 further highlight the robustness of MERLIN in video retrieval tasks.
+The higher $\alpha$ could mitigate the query drift.
 As mentioned in Section 3.4, our assumption is mitigating query drift would preserve the thematic integrity of the initial query by assigning high $\alpha$ value, prioritizing the query and earlier answer embeddings over the most recent answers.
@@
 Furthermore, the average ranking of the target video deteriorates in later rounds, suggesting the query representation has deviated from the user’s original intent.
-Specifically, for MSR-VTT, MERLIN got 44.4/67.60/76.20 for recall@1/5/10 at round 00 respectively but ended up with 61.6/81.20/87.00 respectively at round $5$.
-For MSVD, MERLIN got 52.39/77.16/84.78 for recall@1/5/10 at round 00 respectively but ended up with 56.87/78.51/84.63 respectively at round $5$.
+Specifically, for MSR-VTT, MERLIN got $44.4$/$67.60$/$76.20$ for R@1/5/10 at round 00 respectively but ended up with $61.6$/$81.20$/$87.00$ respectively at round $5$.
+For MSVD, MERLIN got $52.39$/$77.16$/$84.78$ for R@1/5/10 at round 00 respectively but ended up with $56.87$/$78.51$/$84.63$ respectively at round $5$.
+6 Case Study
+The main objective of MERLIN is to improve the ranking of failure cases where the target video is not among the top-ranked candidates.
+At the same time, it is important to keep the success case to stay in the top-ranked candidates while MERLIN proceeds to chat with the user.
+Retrieving the target video among the top-ranked candidates indicates that MERLIN consistently reflects user intention during the conversation.
+To qualitatively verify that MERLIN performs its tasks according to the aforementioned objectives, we reviewed several case studies.
+We focused on how MERLIN brings the rank of failure cases.
+Case study for ActivityNet
+As shown in Figure 4, the initial ranking of the target video was $224$ using a paired query from the dataset.
+However as MERLIN augmented the query using the user’s response, the rank boosted to $36\rightarrow 14\rightarrow 4\rightarrow 1$ as the round proceeded.
+During the conversation, MERLIN was able to understand that the user was looking for a video about Christmas themes, featuring two people, and involving gift wrapping.
+It managed to rank the target video on top with the augmented information.
+Case study for MSVD
+As shown in Figure 5, the initial ranking of the target video was $154$ using a paired query from the dataset.
+However as MERLIN augmented the query using the user’s response, the rank boosted to $14\rightarrow 1\rightarrow 1\rightarrow 1$ as the round proceeded.
+During the conversation, MERLIN was able to understand that the user was looking for a video about the NBA All-Star game, broadcasted on TNT and the scoreboard telling 74:75.
+It managed to rank the target video on top with the augmented information at an early round and managed to keep the top rank during multiple rounds.
+Case study for MSR-VTT
+As shown in Figure 6, the initial ranking of the target video was $361$ using a paired query from the dataset.
+However as MERLIN augmented the query using the user’s response, the rank boosted to $197\rightarrow 14\rightarrow 1\rightarrow 1$ as the round proceeded.
+During the conversation, MERLIN was able to understand the detailed features and gestures of humans featured on “video in mind”.
+It managed to rank the target video on top with the augmented information at an early round and managed to keep the top rank during multiple rounds.
 7 Conclusion
@@
 While our results are promising, we acknowledge that we cannot provide a comprehensive guide for adapting MERLIN to different settings, as we have not extensively explored the impact of changing various components.
-However, the core principle of integrating user feedback to iteratively refine the query embedding appears to be a robust approach, regardless of pipeline components, the specific domain or data modality.
+However, the core principle of integrating user feedback to iteratively refine the query embedding appears to be a robust approach, regardless of pipeline components, the specific domain, or data modality.
 Future work could investigate the generalization of MERLIN to other multimedia retrieval tasks and explore the optimal configurations for different scenarios.
-Other limitation of our approach lies in the use of a human-simulating LLM agent for answering questions based on static video frames.
+Another limitation of our approach lies in the use of a human-simulating LLM agent for answering questions based on static video frames.
 While this agent aims to mimic the human feedback process, it lacks the capability to grasp temporal information and attributes that require a high-level understanding of motion and dynamics.
@@
 This creates a kind of chicken-and-egg problem, where video encoders can benefit from temporal-rich information only when they can understand temporal information effectively.
-Conversely, even if the video question-answering module (or similar counterparts) can handle temporal-rich information, if the video encoder does not possess the same capability, it may not benefit from this information.
+Conversely, even if the video question answering module (or similar counterparts) can handle temporal-rich information, if the video encoder does not possess the same capability, it may not benefit from this information.
 This temporal modeling challenge is a prevalent issue that the community needs to address collectively.
 Acknowledgements
+Nojun Kwak was supported by NRF grant (2021R1A2C3006659) and IITP grants (RS-2022-II220320, RS-2021-II211343), all funded by MSIT of the Korean Government.
 This work was partially supported by Deep Learning Research Group AttentionX.
-Appendix A Algorithm
-Algorithm 1 Iterative video reranking with question answering rounds
-1:encoder $f_{\text{enc}}()$, user query $q\in\mathcal{Q}$, video $v\in\mathcal{V}$, total question-answer round $R$, retrieved top-$k$ videos at round $r$ $\hat{v}^{r}$, $i$-th candidate among top-$k$ videos at round $r$ $\hat{v}^{r}_{i}$
-2:Encode $\mathbf{e}_{q}=f_{\text{enc}}(q)$ given user query $q$
-3:Encode $\mathbf{e}_{v}=f_{\text{enc}}(v)$ given video $v$
-4:Retrieve $\hat{v}^{0}=\textsc{top-k}_{v\in\mathcal{V}}\big{(}\textsc{sim}(\mathbf{e}_{q},\mathbf{e}_{v})\big{)}$ (Equation 1)
-5:Initialize message list $m=[]$
-6:for $r=1$ to $R$ do
-7:     Append metadata of $\hat{v}^{r-1}_{0}$ to $m$
-8:     Generate question $\hat{q}^{r}=\mathcal{M}_{\mathsf{question}}(m)$ (Equation 2)
-9:     Append $\hat{q}^{r}$ to $m$
-10:     Generate frame-level answers $\big{[}\hat{a}^{(r,0)},\dots,\hat{a}^{(r,N)}\big{]}=\mathcal{M}_{\mathsf{answer}}(\hat{q}^{r}:\text{video in mind})$ (Equation 3)
-11:     Aggregate frame-level answers $\tilde{a}^{r}=\mathcal{M}_{\mathsf{aggr}}([\hat{a}^{(r,0)},\dots,\hat{a}^{(r,N)}\big{]})$ (Equation 4)
-12:     Encode $\mathbf{e}_{A^{r}}=f_{\text{enc}}(A^{r})$
-13:     Refine embedding $e=Refine(e_{q},\cdots,e_{A^{r}})$
-14:     Retrieve $\hat{v}^{r}=\textsc{top-k}_{v\in\mathcal{V}}\big{(}\textsc{sim}(\mathbf{e},\mathbf{e}_{v})\big{)}$ (Equation 1)
-15:end for
-16:return Reranked retrieved videos $\hat{v}^{r}_{k}$
-Appendix B Prompt Template
-B.1 Prompt for Question Generation Module
-In order to get useful information from user, it is critical to ask good question that could draw user intention.
-As depicted in Table 2, we set top $1$ ranked video as anchor video and prompted GPT-4o to refer to anchor video’s metadata.
+Appendix A Prompt Template
+A.1 Prompt for Question Generation Module
+To get useful information from a user, it is critical to ask good questions that could elicit the user’s intention.
+As depicted in Table 3, we set the top $1$ ranked video as the anchor video and prompted GPT-4o to refer to the anchor video’s metadata.
 In our case, we used the video’s caption as metadata.
 However, we believe that questions could be more diverse if we could use other data such as Automatic Speech Recognition (ASR) captions, the characteristics of the video, and so on.
-As MERLIN proceeds with the chat with the user (a user-simulating agent), we stacked previous questions and answers and To encourage GPT-4 to generate diverse questions without repeating previous ones..
-B.2 Prompts for Human-Simulating Agent
-As human-simulating agent has two steps for answering to the question regarding “video in mind”, we have two different settings for each step.
-This method lacks in understanding direction, speed and other temporal knowledge as we dealt in the Limitation.
-However, we experimentally showed that our human-simulating agent is helpful in enriching information.
-B.2.1 Prompt for Question-Answering Module
-As depicted in Table 3, we sampled frame from video for every 1 second.
-Then we asynchronously inputed sampled frame and question from MERLIN.
-We prompted GPT-4o to answer in detail about fact and not just answer in “Yes” or “No”.
-However this question-answering module is the part that takes up large portion of API cost so video may be sample in wider stride to lower the API cost.
-B.2.2 Prompt for Aggregation Module
-As depicted in Table 4, we aggregate all the answers generated from the question-answering module.
-We prompted GPT-4o to aggregate multiple answers made with multiple frame at question-answering module and appended an aggregating example.
-Appendix C Case Studies
-The main objective of MERLIN is to improve the ranking of failure cases where the target video is not among the top-ranked candidates.
-At the same time, it is important to keep the success case to stay in the top-ranked candidates while MERLIN proceeds chatting with user.
-Retrieving the target video among the top-ranked candidates indicates that MERLIN consistently reflects user intention during the conversation.
-To qualitatively verify that MERLIN performs its tasks according to the aforementioned objectives, we reviewed several case studies.
-We focused on how MERLIN bring the rank of failure cases.
-C.1 Case study for ActivityNet
-As shown in the Figure 5, initial ranking of target video was $224$ using paired query from dataset.
-However as MERLIN augmented query using user’s response, rank boosted to $36\rightarrow 14\rightarrow 4\rightarrow 1$ as round proceeds.
-During the conversation, MERLIN was able to understand that the user was looking for a video about Christmas themes, featuring two people, and involving gift wrapping.
-It managed to rank the target video on top with augmented information.
-C.2 Case study for MSVD
-As shown in Figure 6, initial ranking of target video was $154$ using paired query from dataset.
-However as MERLIN augmented query using user’s response, rank boosted to $14\rightarrow 1\rightarrow 1\rightarrow 1$ as round proceeds.
-During the conversation, MERLIN was able to understand that the user was looking for a video about NBA All-Star game, broadcasted on TNT and score board telling 74:75.
-It managed to rank the target video on top with augmented information at early round and manage to keep top rank during multiple rounds.
-C.3 Case study for MSR-VTT
-As shown in Figure 7, initial ranking of target video was $361$ using paired query from dataset.
-However as MERLIN augmented query using user’s response, rank boosted to $197\rightarrow 14\rightarrow 1\rightarrow 1$ as round proceeds.
-During the conversation, MERLIN was able to understand that the detailed feature and gesture of human featured on “video in mind”.
-It managed to rank the target video on top with augmented information at early round and manage to keep top rank during multiple rounds.
+As MERLIN proceeds with the chat with the user (a user-simulating agent), we stacked previous questions and answers and encouraged GPT-4 to generate diverse questions without repeating previous ones.
+A.2 Prompts for Human-Simulating Agent
+As a human-simulating agent has two steps for answering the question regarding “video in mind”, we have two different settings for each step.
+This method lacks in understanding direction, speed, and other temporal knowledge as we discussed in Limitation.
+However, we experimentally showed that our human-simulating agent helps enrich information.
+A.2.1 Prompt for Question Answering Module
+As depicted in Table 4, we sampled frames from a video for every 1 second.
+Then we asynchronously input the sampled frames and the question from MERLIN.
+We prompted GPT-4o to answer in detail about facts and not just answer with “Yes” or “No”.
+However, this question answering module is the part that takes up a large portion of API cost so the video may be sampled in a wider stride to lower the API cost.
+A.2.2 Prompt for Aggregation Module
+As depicted in Table 5, we aggregate all the answers generated from the question answering module.
+We prompted GPT-4o to aggregate multiple answers made with multiple frames at the question answering module and appended an aggregating example.
 Details about question generation module in MERLIN
 System prompt
-You are given caption about certain video(anchor video) and query used to retrieve the anchor video. However this video may not be the exact video the I am looking for.
-Your role is to ask question about the video I have in mind to get more information about video. You have 5 rounds and you can only ask one question at a time.
-Focus on attributes like number of people, color, shape and etc.
+You are given a caption about a certain video(anchor video) and a query used to retrieve the anchor video. However, this video may not be the exact video that I am looking for.
+Your role is to ask questions about the video I have in mind to get more information about the video. You have 5 rounds and you can only ask one question at a time.
+Focus on attributes like the number of people, color, shape etc.
 Initial prompt
-This is caption of retrieved video.
-Read the video captions and ask some question to gain more information to help find out exact video.
-Some video may not have caption due to API error saying sorry I can’t provide blah blah.
+This is the caption of the retrieved video.
+Read the video captions and ask some questions to gain more information to help find out the exact video.
+Some videos may not have a caption due to an API error saying sorry I can’t provide blah blah.
 Captions for video: {anchor video’s caption}
 Question:
-Question-answering round prompt
+Question answering round prompt
 answer: {Aggregated answer from user-simulating Agent}
-Based on answer, here’s caption of reranked video.
+Based on the answer, here’s the caption of the reranked video.
 caption: {reranked top1 video caption as anchor caption}
@@
 - 0.75
-Table 2: The instruction and specification for question generation module in MERLIN using GPT-4o.
-After initial retrieval at round 00, MERLIN generates question with initial prompt using information of anchor video’s caption.
-After user answers to the question, MERLIN rerank the and generate question using new anchor and question-answering round prompt.
-Details about human-simulating agent (question-answering module)
+Table 3: The instruction and specification for the question generation module in MERLIN using GPT-4o.
+After initial retrieval at round 00, MERLIN generates a question with an initial prompt using the information of the anchor video’s caption.
+After the user answers the question, MERLIN reranks the and generates a question using a new anchor and question answering round prompt.
+Details about human-simulating agent (question answering module)
 System Prompt
-You are a helpful assistant that answer the question with details. Don’t jsut answer in yes or no. Provide more details(about facts) about the image that might help the questioner.
+You are a helpful assistant that answers the question with details. Don’t just answer in yes or no. Provide more details(about facts) about the image that might help the questioner.
 Input format
@@
 - 1 second
-Table 3: The instruction and specification for video question-answering human-simulating agent using GPT-4o (question-answering module).
-Details about human-simulating agent (aggregation module)
+Table 4: The instruction and specification for video question answering human-simulating agent using GPT-4o (question answering module).
+Details about the human-simulating agent (aggregation module)
 System Prompt
@@
 Your role is to review all of the individual answers and summarize them to provide a final answer to the original question.
-When making final answer, don’t user unnecessary words like ‘Based on the individual answers provided by the VQA model,’.
-Just answer to the question.
+When making the final answer, don’t use unnecessary words like ‘Based on the individual answers provided by the VQA model,’.
+Just answer the question.
 For example, if the question is “Did a cookie appear in the video?” and the individual answers from the frames are [“No”, “No”, “Yes”, “No”], then since a cookie appeared in the 3rd frame, you should summarize and answer the question as “Yes”.
-Length of aggregated answer should be around 30~35 words.
+The length of the aggregated answer should be around 30~35 words.
 Input format
 Question: {Question from MERLIN}
-VQA Answer: {Answers from question-answering module}
+VQA Answer: {Answers from question answering module}
 Aggregated Answer:
@@
 - 0.5
-Table 4: The instruction and specification for video question-answering human-simulating agent using GPT-4o (aggregation module).
+Table 5: The instruction and specification for video question answering human-simulating agent using GPT-4o (aggregation module).
 Figure 4: Qualitative evaluation of MERLIN on ActivityNet. sample: v_juiMCvZUYwk.